Find zα/2 for each of the following confidence levels used in estimating the population proportion.

Round your answers to 3 decimal places.)

zα/2
a. 90%
b. 98%
c. 85%
d. 95%
e. 99%
Answer #1

Solution :- given that, is a) At 90% confidence level the z 2:1 - 90% <= 1-0.90 soil 42 = 0.05 242 = 20.05 = 1.645 24 - 1.645Date is d. At 95% confidence level the 2 Kel-95 x= 1-0.95 x = 0.05 42 = 0,025 Z42 = 20.025 = 1.960 242= 1.960 is At 99% confi